Circular No. 7274

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1859+226 A. Wood and D. A. Smith, Massachusetts Institute of Technology

(MIT); and F. E. Marshall and J. Swank, Goddard Space Flight Center (GSFC), on behalf of the RXTE All-Sky Monitor (ASM) team at MIT and GSFC, report the emergence of a new x-ray transient: “The source’s intensity was measured to rise slowly during a ∼ 12-hr interval to 160 ± 15 mCrab (2–12 keV) on Oct. 9.658 UT and 250 mCrab by Oct. 10.523, confirmed by a PCA observation (with source position given by Markwardt et al., below). The rise was consistent with a linear increase at a rate of ∼ 6 mCrab/hr. Extrapolation indicates that the outburst began around Oct. 8.55. The ASM spectrum is roughly Crab-like and shows no indication of evolution. Neither the ASM catalogue nor SIMBAD lists any known x-ray sources in or near the positional error box. Observations at other wavelengths would be timely.”

C. B. Markwardt, University of Maryland and GSFC; and F. E. Marshall and J. H. Swank, GSFC, write: “The RXTE PCA observed the transient source XTE J1859+226 on Oct. 10.56–10.60 UT for ∼ 2000 s. The intensity of the source was 1.7 × 10^{−8} erg s^{−1} cm^{−2} in the band 2–60 keV (= 250 mCrab in the band 2–10 keV), with fractional r.m.s. variability of ∼ 32 percent. PCA scans during the observation determined a source position of α = 18^{h}58^{m}35^{s}, δ = +22^{o}39′.3 (equinox 2000.0), with a 95-percent uncertainty radius of 1′, which also accounts for the intrinsic variability of the source. A power spectrum shows a quasiperiodic oscillation and two harmonics, the fundamental being centered at 0.45 Hz (FWHM 0.11 Hz), with a fractional r.m.s. amplitude of ∼ 17 percent. No high-frequency pulsations or variability were seen. The energy spectrum can be modeled as a power law with photon index 1.7 and neutral hydrogen absorption of ∼ 3 × 10^{21} cm^{−2}. These timing and spectral characteristics together are reminiscent of a galactic blackhole candidate. Additional PCA observations are planned.”

COMET C/1999 S3 (LINEAR) Total visual magnitude estimates: Sept. 27.88 UT, 12.6 (K. Hornoch,

Lelekovice, Czech Republic, 0.35-m reflector; moonlight); 29.82, 12.5 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); Oct. 1.79, 12.7 (W. Hasubick, Buchloe, Germany, 0.44-m reflector); 2.82, 12.5 (M. Lehký, Hradec Králové, Czech Republic, 0.42-m reflector); 2.92, 12.3 (G. Comello, Roden, The Netherlands, 0.31-m reflector); 6.10, 12.6 (Bouma).

1999 October 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green